OpenStudy (amistre64):

integrate: cot^-1(x) dx

OpenStudy (anonymous):

u sure it is integrable?

OpenStudy (amistre64):

yep; just a pain in the butt to do i believe :)

OpenStudy (anonymous):

u sure it is integrable?

OpenStudy (anonymous):

guess so

OpenStudy (amistre64):

int by parts using the cot-1 as the deriving part; and dx is the integrating part

OpenStudy (amistre64):

.... man these are tricky lol

OpenStudy (amistre64):

\begin{array}c &&dx&\\ +&cot^{-1}(x)&x\\ -&Dx(cot^{-1}(x))&x^2/2 \end{array}

OpenStudy (amistre64):

\[cot^{-1}(x)=y\] \[x=cot(y)\] \[1=-csc^2(y)dy/dx\] \[-\frac{1}{csc^2(y)}=y'\]

OpenStudy (amistre64):



OpenStudy (amistre64):

csc(y) = sqrt(x^2+1) csc^2(y) = x^2 + 1

OpenStudy (anonymous):

is the question: \[\int\limits_{}^{}\cot^{-1} xdx\]

OpenStudy (amistre64):

\[Dx(cot^{-1}(x))=-\frac{1}{x^2+1}\]

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

cot^-1(x) = sinx/cosx?

OpenStudy (amistre64):

not quite, the inverse cotangent function doesnt equation the tangent function :)

OpenStudy (amistre64):

doesnt *equal* the .....

OpenStudy (amistre64):

\[[uv]'=u\ dv+v\ du\] \[uv =\int u\ dv+\int v\ du\] \[ \int u\ dv\ =\ uv\ -\int v\ du\] \[\int cot^{-1}(x)\ dx\] \[u=cot^{-1}(x);\ dv = dx\] \[du=-\frac{1}{x^2+1}dx; \ v = x\] \[ \int cot^{-1}(x)\ dx\ =\ cot^{-1}(x)\ x\ +\int \frac{x}{x^2+1}dx\] \[=x\ cot^{-1}(x) +\frac{1}{2}\ ln(x^2+1)+C\] right?

OpenStudy (anonymous):

yeah that's good!